Show that it is not possible for a photon to be completely absorbed by a free electron.

Grade:10

1 Answers

Aditi Chauhan
9 years ago
Sol. If the I undergoes an elastic collision with a photon. Hen applying energy conservation to this colllison We get, hC/λ + m base 0 c^2 = mc^2 And applying conservation of momentum h/λ Mass of e = m = m base 0/√1 – v^2 /C^2 From above equation it can be easily shown that V = C or V = 0 Both of these results have no physical meaning hence it is not possible for a photon to e completely absorbed ny a freed electron.